OpenStudy (anonymous):

Polar Graphs r=2+3sin(theta) Find the slope of the tangent line at (2, pi)

OpenStudy (anonymous):

dy/dx = (r'(theta)*sin(theta) + r(theta)cos(theta)) / (r'(theta)*cos(theta) - r(theta)sin(theta)) sin(pi) = 0 cos(pi) = -1 r'(theta) = 3cos(theta) r' (pi) = 3cos(pi) = -3 dy/dx = (-3 * 0 + 2*(-1)) / (-3 * (-1) - 2 * 0) dy/dx = -2/3

OpenStudy (anonymous):

oh!!! I missed the 2+3sintheta in my r(theta) so i still had a 0 on top! Thank you so much......teaching myself Calc BC is kinda frustrating =(

OpenStudy (anonymous):

to be honest i dont like this subject(polar graphs)